Dry air is about 20.95% oxygen by volume. Assuming STP, how many oxygen molecules are in a 75.0 g sample of air? The density of air is 1.19 g/L.
1
Expert's answer
2020-01-20T08:49:31-0500


At STP temperature, T=273.15K, pressure, P=101325Pa.

mass, m=75g and density, d=1.19g/l, R=8.315

number of moles,"n=PV\/RT"

"V=m\/d"

V=75/1.19= 63cm3

n=(101325*63*10-6)/(8.315*273.15)

n=2.81*10-3moles